The line that contains (-2,3) and (6,b) has a slope m = -4. Find the value of b.
Answer by Cromlix(4381)

The gradient = b - 3/6 -(-2) = b - 3/ 8
b - 3/8 = -4
Cross multiply
b - 3 = -32
b = -29
-29 - 3/8 => -32/8 => -4